# Past, Present, and Future of Bug Tracking in the Generative AI Era

### v1

@@ 2.1.5. Software as a service (SaaS), DevOps and Automation Era (2010s-2022)
Era Bug Reporting Bug Reproduction Bug Fixing Verification Bug Classification Support Representative Tools What is New? Early Digital Era (1940s-1970s) Bugs were written to the notebooks, sticky notes, spreadsheets, etc. There are a certain number of testing methods in the system, and with these testing methods, the bug is reproduced. For notational bugs, bugs are fixed by correcting the wrong place in the notebook. For the concurrency bugs, the bugs are fixed by using a dummy program to adjust the signals. The same methodology as the Bug Reproduction. If all tests are passed, then the bug is verified. Bug types or severity are not consistently captured Paper Logs, Index Cards, Notebooks NA Pre-Internet Era (1970s-1980s) Reported via email/phone or customer support. Software shipped on floppy disks with registration forms; users mailed forms and faulty disks back to developers. Customer support receives bug reproduction steps from the user via email or phone call, then writes these steps to the text file. Then these steps are passed to the developer to reproduce the bug. Manual code changes, shared via floppy disk or a local network. Manual testing, often with a documented checklist. Simple severity labels, often inconsistently used Email, text editors, simple databases (e.g., dBase), network file shares, Lotus 1-2-3.242424https://en.wikipedia.org/wiki/Lotus_1-2-3 Introduction of remote communication for bug reporting (email, phone) and use of digital storage for transferring software and fixes. Internet Era (1980s–1990s) Reports were logged into the system by a user with some structure. Detailed Steps to Reproduce (S2R) were logged in the system, accessible to the whole team. Code fixed and committed to VCS. Updates were often manual in the bug report. Formal QA teams re-tested fixes; verification was largely manual and sequential. Simple severity labels, often inconsistently used GNATS, Debuggs, CMVC Structured digital bug databases introduced, enabling team-wide visibility and linking bugs to code changes. Web-Based Era (2000s) Web UI support, bug reports were treated as work items in a sprint Detailed S2R were logged in the system, accessible to the whole team. Fixes tied directly to modern VCSs (Git, etc.). Systems started auto-linking commits/pull requests to bug reports. QA remained formalized but incorporated automation (test scripts, regression suites) and parallel workflows. Manual categorization with some automation Bugzilla, MantisBT, Trac, Redmine, Jira, Microsoft Team Foundation Server (TFS) Shift to web-based collaboration, integration with Agile workflows, and more formalized bug management processes. SaaS, DevOps and Automation Era (2010s–2022) Web-based, collaborative platforms integrated with other development tools. Logs/screenshots attached to tickets; testing environment integration became standard Fixes were part of a CI/CD pipeline. Fixes were tracked via VCSs. Automated testing became a standard practice, with test results often linked directly to the bug ticket. Tagging, priority queues, category rules GitHub Issues, GitLab Issues, YouTrack, Azure DevOps Fully integrated bug tracking into DevOps toolchains with CI/CD, automated testing, and real-time collaboration.
2.2. Current System
@@
Figure 1 presents a typical workflow of the traditional (SaaS, DevOps, and Automation Era) bug tracking system, showing the interactions among different roles. The process starts when a user creates a bug report. Customer support then attempts manual bug reproduction. If reproduction is unsuccessful, additional details are requested from the user, and the attempt is repeated until the bug can be reliably reproduced. Once reproduction is successful, the bug proceeds to classification. After classification, a bug–feature traceability step is performed, followed by the validity check. In the traditional literature, classification and assignment together are often referred to as “bug triage,” but in our discussion, we separate these into two distinct stages: classification and assignment.
@@
Figure 1. Overview of Traditional Bug Tracking System Workflow:
2.3. Challenges of Current System
Given the traditional bug-tracking process described above, it is important to examine the key challenges that hinder its effectiveness. While the primary objective is to identify and resolve bugs within an intended time span, the whole process, which starts with bug reporting and ends with final deployment and user verification, faces many obstacles. These issues are not just caused by technical reasons; they often arise from human factors (Sun, 2011; Zimmermann et al., 2010). In the following, we have summarized the key challenges.
@@ 2.3.1. Challenges of Initial Bug Report Creation
The bug-fixing process is fundamentally tied to the quality of the information provided in the bug report, which is one of the cornerstones of a bug tracking system. A primary challenge is that information in these reports is often incomplete or vague (Sun, 2011; Zimmermann et al., 2010; Breu et al., 2009). This can lead to reports that lack clear reproduction steps and crucial environmental details. Missing environment variables, such as product version and operating system (OS) version in a report, and missing severity are two examples of bug tracking process smells (Qamar et al., 2022), specifically about bug reports. As a result of these bad practices, customer support staff and developers are forced to engage in a time-consuming ”ping-pong” of communication with the reporter, requesting additional information (Qamar et al., 2022). This back-and-forth exchange significantly slows down the bug resolution process and increases the number of unresolved issues (Zimmermann et al., 2009). The problem is often compounded by the passive nature of traditional bug tracking systems, which are little more than interfaces for relational databases and offer minimal support for reporters to provide the necessary context (Zimmermann et al., 2009). In addition, reporters who are not technically experienced may struggle to articulate problems in a structured way or find bug tracking interfaces difficult to use, further reducing the clarity and usefulness of bug reports.
@@ 2.3.2. Challenges of Bug Reproduction
Once a bug is reported, the next step is consistently reproducing it in a controlled environment. This task can be surprisingly difficult due to several reasons, such as discrepancies between the original user environment and the development environment or lack of information in the bug report (Li et al., [n. d.]; Vyas et al., 2014). Bugs that are not raised during obvious conditions can increase the challenge of reproducibility. There are bugs called Heisenbugs (Gray, 1986) that change their behavior or disappear entirely when debugging or observing them, making them difficult to reproduce (Huang and Zhang, 2016). Bugs that are caused by concurrency issues can be given as an example to this class. The rate of non-reproducible bugs is far from negligible; in their study, Goyal et al. (Goyal and Sardana, 2019) mined bug reports from four large-scale open source software projects of the Bugzilla repository and found that non-reproducible bugs range from 12.77% to 24.26% of all reported bugs. Failing to reliably reproduce an issue often leads to frustration and wasted effort for both the development team and users.
@@ 2.3.3. Challenges of Bug Classification
Manual classification is labor-intensive and time-consuming, which involves a series of sequential actions, demands substantial problem-solving abilities and careful consideration. The large volume of incoming bug reports can easily overwhelm teams, leading to a mounting backlog of issues and significant delays in the software development lifecycle (Kukkar et al., 2020). Reports often contain incomplete information and ambiguous descriptions (Bettenburg et al., 2008), which may complicate accurate classification.
@@ 2.3.4. Challenges of Traceability of Bugs to Other Artifacts
Traceability of the bug is the ability to trace it in the whole bug lifecycle using artifact links. Examples for these artifacts can be bug fixing commits (Rath et al., 2018) and Pull Request (PR) of the bug fix (Yaşa et al., 2025). Due to a lack of set rules and oversight, developers often fail to link their PRs to the specific issues they are meant to solve (Ruan et al., 2019). This disconnect makes it difficult to understand why certain code was changed, leading to a loss of valuable information that could be useful when problems arise. Another important aspect is Requirements Traceability (RT), which refers to the ability to track and follow the life of a requirement across the project lifecycle (Gotel and Finkelstein, 1994). RT links are valuable during maintenance because they allow developers to quickly understand the impact of changes, trace defects back to their originating requirements, and ensure that modifications remain consistent with stakeholder needs. However, creating and maintaining these links is laborious, as it often requires manual effort to analyze requirements, design artifacts, code, and test cases, and then establish explicit relationships between them. Lack of traceability can raise problems in the future of that software project, such as the reduced effect of change impact analysis (Aung et al., 2020). Automated techniques to construct these links would save resources.
@@ 2.3.5. Challenges of Detecting Invalid Bugs
Another possible challenge is to detect if the bug is valid or not. Some of the causes of invalid bugs are insufficient background knowledge, misunderstanding of functionality, misunderstanding of environment, error in testing, and error in the external system, according to Sun et al. (Sun, 2011).
While automated bug classification of valid/invalid bugs using ML has shown promise in reducing manual effort, its adoption faces several practical challenges. A major issue is the quality and imbalance of real-world bug datasets. Projects often contain a large proportion of non-bug or duplicate reports compared to true defects, which biases models toward the majority classes and reduces performance on the more critical minority classes. Beyond imbalance, bug reports themselves are highly heterogeneous: they may be written in natural language with varying levels of detail, include incomplete or noisy information, or use project-specific jargon that makes feature extraction difficult. Additionally, labeling ground truth data is labor-intensive and prone to human inconsistency, further limiting the reliability of training data. These challenges collectively highlight why practical deployment of automated bug classification remains non-trivial despite promising research results.
@@ 2.3.6. Challenges of Bug Localization
A primary challenge in debugging is the initial task of identifying the root cause of an issue. This can be a challenging task, especially in large and complex codebases where developers must sift through a large number of potential culprits to find responsible lines of code (Ghosh and Singh, 2019). The difficulty is compounded in modern environments by the distributed and interconnected nature of applications, where issues may traverse multiple system layers and require expertise with a range of specialized tools (Ghosh and Singh, 2019).
2.3.7. Challenges of Creating Bug Fix
While a bug might be identified and its root cause understood, the actual process of changing the code is far from trivial. Typically, a change made to a software system may result in an undesirable side effect or ripple effect to the rest of the system (Arnold, 1996). There is a unique set of challenges that can impact code quality, introduce new defects, code smells, and increase the overall cost of development. Changed code might affect other parts of the code, which is challenging to detect, as change impact analysis is being used to measure the side effects of code changes using several metrics (Göçmen et al., 2025). Lack of documentation might cause challenges in the code change process as well. Without clear documentation, developers may not understand why a particular piece of code was written a certain way.
@@ 2.3.8. Challenges of Bug Fix Verification
Verifying a bug fix is the final, critical step in the bug-fixing lifecycle. It is the process of confirming that the bug is truly resolved and that the fix has not introduced any new problems. The primary technical challenge in bug fix verification is the risk of regression. A bug fix, especially in a complex system, can inadvertently break existing, previously functional code. This requires testers to go beyond merely checking if the original bug is fixed and perform extensive regression testing, which can be time-consuming and difficult without robust automated test suites. Another technical challenge in this step is reproducibility, which is also a challenge mentioned in Section 2.3.2. The scope of testing is also a challenge, as a tester must correctly identify all the areas of the application that could have been affected by the fix, a task that becomes exponentially more difficult with tightly coupled code and a lack of clear documentation. Beyond the technical aspects, bug fix verification is significantly affected by human and organizational factors. Deak et al. (Deak et al., 2016) show that testers are affected by negative factors categorized as: lack of influence and recognition, unhappy with management, technical issues, lack of organization, time pressure, boredom, poor relationships with developers, and working environment issues.
@@ 2.3.9. Challenges of Asynchronous Communication
In modern bug tracking workflows, much of the interaction between stakeholders takes place through asynchronous communication that is, exchanges like emails, issue comments, or chat messages that do not require participants to be online at the same time. The primary challenge of using asynchronous communication is the inherent lack of real-time context and immediate feedback. Unlike synchronous methods such as live meetings or screen shares, asynchronous bug reporting and resolution can lead to significant delays (Eren et al., 2023). A developer may need to ask for more information about the bug, such as specific S2R for it or screenshots from a different browser, but the response may not come for hours or even days. This time lag, often referred to as communication latency or communication gap, can stall the entire bug-fixing process and cause a loss of momentum (Eren et al., 2023). It also makes it difficult to have a nuanced, back-and-forth discussion, which is often necessary to fully understand and diagnose a complex issue. Study (Korkala et al., 2006) shows that when the reliance on asynchronous tools increases, the software defect rates increase accordingly.
@@ 3. AI-Powered Bug Tracking Framework
To mitigate the problems of the current bug tracking systems that are detailed in Section 2, we propose a visionary bug tracking framework leveraging LLM/AI agents. We address some of the limitations highlighted in the previous section, from the subjectivity and inefficiency of manual bug reporting and classification to the complexities of bug localization and the risk of regression during fix verification. By integrating LLM/AI agents into bug tracking, we aim to enable automation and intelligence at multiple stages of bug tracking. LLMs’ role begins with assisting in bug report creation and enhancement and continues through reproduction, classification, traceability link creation, validation, bug assignment, localization, fixing, verification, and deployment.

### v2

@@ 2.1.5. DevOps and Automation Era (2010s-2022)
Era Bug Reporting Bug Reproduction Bug Fixing Verification Bug Classification Support Representative Tools What is New? Early Digital Era (1940s-1970s) Bugs were written to the notebooks, sticky notes, spreadsheets, etc. There are a certain number of testing methods in the system, and with these testing methods, the bug is reproduced. For notational bugs, bugs are fixed by correcting the wrong place in the notebook. For the concurrency bugs, the bugs are fixed by using a dummy program to adjust the signals. The same methodology as the Bug Reproduction. If all tests are passed, then the bug is verified. Bug types or severity are not consistently captured Paper Logs, Index Cards, Notebooks Problems that may occur in hardware and software solutions should be addressed. Pre-Internet Era (1970s-1980s) Reported via email/phone or customer support. Software shipped on floppy disks with registration forms; users mailed forms and faulty disks back to developers. Customer support receives bug reproduction steps from the user via email or phone call, then writes these steps to the text file. Then these steps are passed to the developer to reproduce the bug. Manual code changes, shared via floppy disk or a local network. Manual testing, often with a documented checklist. Simple severity labels, often inconsistently used Email, text editors, simple databases (e.g., dBase), network file shares, Lotus 1-2-3.242424https://en.wikipedia.org/wiki/Lotus_1-2-3 Introduction of remote communication for bug reporting (email, phone) and use of digital storage for transferring software and fixes. Internet Era (1980s–1990s) Reports were logged into the system by a user with some predefined structure. Detailed Steps to Reproduce (S2R) were logged in the system, accessible to the whole team. Code fixed and committed to VCS. Updates were often manual in the bug report. Formal QA teams re-tested fixes; verification was largely manual and sequential. Simple severity labels, often inconsistently used GNATS, Debuggs, CMVC Structured digital bug databases introduced, enabling team-wide visibility and linking bugs to code changes. Web-Based Era (2000s) Web UI support, bug reports were treated as work items in a sprint Detailed S2R were logged in the system, accessible to the whole team. Fixes tied directly to modern VCSs. Systems started auto-linking commits/PRs to bug reports. QA remained formalized but incorporated automation (test scripts, regression suites) and parallel workflows. Manual categorization with some automation Bugzilla, MantisBT, Trac, Redmine, Jira, Microsoft Team Foundation Server (TFS) Shift to web-based collaboration, integration with Agile workflows, and more formalized bug management processes. DevOps and Automation Era (2010s–2022) Web-based, collaborative platforms integrated with other development tools. Logs/screenshots attached to tickets; testing environment integration became standard Fixes were part of a CI/CD pipeline. Fixes were tracked via VCSs. Automated testing became a standard practice, with test results often linked directly to the bug ticket. Tagging, priority queues, category rules GitHub Issues, GitLab Issues, YouTrack, Azure DevOps, Jira Fully integrated bug tracking into DevOps toolchains with CI/CD, automated testing, and real-time collaboration.
2.2. Traditional System
@@
Figure 1 presents a typical workflow of the traditional bug tracking system in DevOps, and Automation Era, showing the interactions among different roles. The process starts when a user creates a bug report. Customer support then attempts manual bug reproduction. If reproduction is unsuccessful, additional details are requested from the user, and the attempt is repeated until the bug can be reliably reproduced. Once reproduction is successful, the bug proceeds to classification. After classification, a bug–feature traceability step is performed, followed by the validity check. In the traditional literature, classification and assignment together are often referred to as “bug triage,” but in our discussion, we separate these into two distinct stages: classification and assignment.
@@
Figure 1. Overview of Traditional Bug Tracking System Workflow
A diagram showing the traditional bug tracking workflow.
2.3. Challenges of Traditional System
Given the traditional bug tracking process described above, it is important to examine the key challenges that hinder its effectiveness. While the primary objective is to identify and resolve bugs within an intended time span, the whole process, which starts with bug reporting and ends with final deployment and user verification, faces many obstacles. These issues are not just caused by technical reasons; they often arise from human factors (Sun, 2011; Zimmermann et al., 2010). In the following, we have summarized the key challenges.
@@ 2.3.1. Challenges of Initial Bug Report Creation
The bug-fixing process is fundamentally tied to the quality of the information provided in the bug report, which is one of the cornerstones of a bug tracking system (Zimmermann et al., 2010). A primary challenge is that information in these reports is often incomplete or vague (Sun, 2011; Breu et al., 2009). This can lead to reports that lack clear reproduction steps and crucial environmental details. As a result of these bad practices, customer support staff and developers are forced to engage in a time-consuming ”ping-pong” of communication with the reporter, requesting additional information (Qamar et al., 2022). This back-and-forth exchange significantly slows down the bug resolution process and increases the number of unresolved issues (Zimmermann et al., 2009). The problem is often compounded by the passive nature of traditional bug tracking systems, which are little more than interfaces for relational databases and offer minimal support for reporters to provide the necessary context (Zimmermann et al., 2009). In addition, reporters who are not technically experienced may struggle to articulate problems in a structured way or find bug tracking interfaces difficult to use, further reducing the clarity and usefulness of bug reports.
@@ 2.3.2. Challenges of Bug Reproduction
Once a bug is reported, the next step is consistently reproducing it in a controlled environment. This task can be rather difficult due to several reasons, such as discrepancies between the original user environment and the development environment or lack of information in the bug report (Li et al., ; Vyas et al., 2014). Software defects that manifest under non-deterministic conditions significantly complicate the process of reproducibility. A classic example is the ’Heisenbug’ (Gray, 1986), a term describing bugs that alter their behavior or disappear entirely when being observed or debugged. Such defects, often exemplified by concurrency issues, remain notoriously difficult to systematically replicate (Huang and Zhang, 2016).
The rate of non-reproducible bugs is far from negligible; in their study, Goyal et al. (Goyal and Sardana, 2019) mined bug reports from four large-scale open source software projects of the Bugzilla repository and found that non-reproducible bugs range from 12.77% to 24.26% of all reported bugs. Failing to reliably reproduce an issue often leads to frustration and wasted effort for both the development team and users.
@@ 2.3.3. Challenges of Bug Classification
Manual classification is labor-intensive and time-consuming, which involves a series of sequential actions, demands substantial problem-solving abilities and careful consideration. The large volume of incoming bug reports can easily overwhelm teams, leading to a mounting backlog of issues and significant delays in the software development lifecycle (Kukkar et al., 2020). The prevalence of incomplete and ambiguous descriptions in bug reports poses a significant challenge to accurate classification (Bettenburg et al., 2008). For instance, Zhang et al. (Zhang et al., 2017) found that 78.1% of Mozilla and Eclipse reports consist of fewer than 100 words, creating a lack of information that prolongs the resolution process.
@@ 2.3.4. Challenges of Traceability of Bugs to Other Artifacts
Traceability of a bug refers to the ability to track it throughout its entire lifecycle using artifact links. Examples for these artifacts can be bug-fixing commits (Rath et al., 2018) and pull requests (PRs) of the bug fix (Yaşa et al., 2025). Due to a lack of set rules and oversight, developers often fail to link their PRs to the specific issues they are meant to solve (Ruan et al., 2019). This disconnect makes it difficult to understand why certain code was changed, resulting in the loss of valuable information that could be useful when problems arise. Another important aspect is Requirements Traceability (RT), which refers to the ability to track and follow the life of a requirement across the project lifecycle (Gotel and Finkelstein, 1994). RT links are valuable during maintenance because they allow developers to quickly understand the impact of changes, trace defects back to their originating requirements, and ensure that modifications remain consistent with stakeholder needs. However, creating and maintaining these links is laborious, as it often requires manual effort to analyze requirements, design artifacts, code, and test cases, and then establish explicit relationships between them. Lack of traceability can raise problems in the future of that software project, such as the reduced effect of change impact analysis (Aung et al., 2020). Automated techniques to construct these links would save resources.
@@ 2.3.5. Challenges of Detecting Invalid Bugs
Another possible challenge is to detect if the bug is valid or not. Some of the causes of invalid bugs are insufficient background knowledge of the reporter, misunderstanding of functionality, misunderstanding of environment, error in testing, and error in the external system, according to Sun et al. (Sun, 2011).
While automated detection of invalid bug reports using ML has shown promise in reducing manual effort, its adoption faces several practical challenges. A major issue is the quality and imbalance of real-world bug datasets. Projects often contain a large proportion of invalid or duplicate reports compared to valid bug reports, which biases models toward the majority classes and reduces performance on the more critical minority classes (Tantithamthavorn et al., 2018). Beyond imbalance, bug reports themselves are highly heterogeneous: they may be written in natural language with varying levels of detail, include incomplete or noisy information, or use project-specific jargon that makes feature extraction difficult. Additionally, labeling ground truth data is labor-intensive and prone to human inconsistency, further limiting the reliability of training data. These challenges collectively highlight why practical deployment of automated bug validity determination remains non-trivial despite promising research results.
@@ 2.3.6. Challenges of Bug Localization
A primary challenge in debugging is the initial task of identifying the root cause of an issue. This task is particularly demanding in complex codebases, where developers must isolate the root cause from a vast number of potential sources (Ghosh and Singh, 2019). The difficulty is compounded in modern environments by the distributed and interconnected nature of applications, where issues may traverse multiple system layers and require expertise with a range of specialized tools (Ghosh and Singh, 2019).
2.3.7. Challenges of Creating a Bug Fix
Addressing invalid bug reports, which might stem from misconfigurations, incompatible data usage, or simple permission issues rather than actual code defects, presents a significant workflow bottleneck. Currently, the resolution of these issues relies heavily on manual intervention by customer support staff, who must diagnose the environmental mismatch and apply no-code fixes such as directing users to make changes in their configuration files or user access levels. This dependency on manual intervention creates substantial latency; even when the solution is as simple as a permission update from the user side, the asynchronous nature of communication for support often forces end-users to wait hours or even days for a resolution. Moreover, because these operational changes occur outside the rigorous CI/CD pipelines used for code patches, they are prone to human error and lack verification and automation.
Regarding valid bug reports, while a bug might be identified and its root cause understood, the actual process of changing the code is far from trivial. Typically, a change made to a software system may result in an undesirable side effect to the rest of the system (Arnold, 1996). There is a unique set of challenges that can impact code quality, introduce new defects and code smells, and increase the overall development cost. Changes to the code might affect other parts of the code, which can be challenging to detect, as change impact analysis is used to measure the side effects of code changes using several metrics (Göçmen et al., 2025). A lack of documentation may also cause challenges in the code change process. Without clear documentation, developers may struggle to understand why a particular piece of code was written in a certain way. Furthermore, the nature of the fault itself can render the fixing process exceptionally difficult. Specifically, challenges arise from Mandelbugs (Grottke and Trivedi, 2005) and their subtype, Heisenbugs. While Mandelbugs are characterized by complex, non-deterministic activation mechanisms, Heisenbugs specifically exhibit behavior that shifts under observation or during debugging. These transient issues, often rooted in race conditions or timing delays, complicate the classification process because they are inherently difficult to isolate, reproduce, and verify.
Beyond the complexity of the fault logic, many failures cannot be fixed through source code modifications or no-code fixes alone. This category of defects originates from the state of the execution environment rather than internal algorithmic errors. For example, a failure might be triggered by resource exhaustion, like a memory leak in a third-party library, or inconsistent schema states between shared databases. In such instances, a standard source-level patch is insufficient; the resolution instead requires environmental change, such as reconfiguring container resource limits.
@@ 2.3.8. Challenges of Bug Fix Verification
Verifying a bug fix is the final, critical step in the bug-fixing lifecycle. It is the process of confirming that the bug is truly resolved and that the fix has not introduced any new problems. The primary technical challenge in bug fix verification is the risk of regression. A bug fix, especially in a complex system, can inadvertently break existing, previously functional code (Yoo and Harman, 2012). This requires testers to go beyond merely checking if the original bug is fixed and perform extensive regression testing, which can be time-consuming and difficult without robust automated test suites. Another technical challenge in this step is reproducibility, which mirrors the difficulties discussed in Section 2.3.2. Specifically, this is closely linked to the challenge of flaky tests, tests that fail intermittently without any code changes (Luo et al., 2014). The scope of testing is also a challenge, as testers must correctly identify all areas of the application that could have been affected by the fix, a task that becomes exponentially more difficult with tightly coupled code and a lack of clear documentation (Arnold, 1996). Beyond the technical aspects, bug fix verification is significantly affected by human and organizational factors. Deak et al. (Deak et al., 2016) demonstrate that testers are affected by negative factors, including a lack of influence and recognition, dissatisfaction with management, technical issues, inadequate organization, time pressure, boredom, poor relationships with developers, and issues with the working environment.
@@ 2.3.9. Challenges of Asynchronous Communication
In modern bug tracking workflows, much of the interaction between stakeholders takes place through asynchronous communication, that is, exchanges like emails, issue comments, or chat messages that do not require participants to be online at the same time. The primary challenge of using asynchronous communication is the inherent lack of real-time context and immediate feedback. Unlike synchronous methods such as live meetings or screen shares, asynchronous bug reporting and resolution can lead to significant delays (Eren et al., 2023). A developer may need to request more information about the bug, such as specific S2R for it or screenshots from a different browser, but the response may not arrive for hours or even days. This time lag, often referred to as communication latency or communication gap, can stall the entire bug-fixing process and cause a loss of momentum (Eren et al., 2023). It also makes it difficult to have a nuanced, back-and-forth discussion, which is often necessary to fully understand and diagnose a complex issue. Korkala’s study (Korkala et al., 2006) shows that when the reliance on asynchronous tools increases, the software defect rates increase accordingly.
@@ 3. AI-Powered Bug Tracking Framework
To mitigate some of the problems of the traditional bug tracking systems detailed in Section 2, we propose a visionary bug tracking framework that leverages LLM/AI agents. We address some of the limitations highlighted in the previous section, including the subjectivity and inefficiency of manual bug reporting and classification, as well as the complexities of bug localization and the risk of regression during fix verification. By integrating LLM/AI agents into bug tracking, we aim to enable automation and intelligence at multiple stages of bug tracking. LLMs’ role begins with assisting in bug report creation and enhancement and continues through reproduction, classification, traceability link creation, validation, bug assignment, localization, fixing, verification, and deployment.